# Chain-of-Instructions: Compositional Instruction Tuning on Large Language Models

### v2

@@ 5 Related Work
Several NLP work have investigated the capability of Transformer model on compositional problems including algorithm and math problems Dziri et al. (2023), compositional semantics Drozdov et al. (2022), and multi-hop question-answering (QA) tasks Trivedi et al. (2022). Dziri et al. (2023) highlight how Transformer models often struggle with compositional mathematics computation or program executions Nye et al. (2022); Saparov and He (2022). Drozdov et al. (2022) introduce a new prompting method which first decomposes the compositional questions or sentences Keysers et al. (2019); Kim and Linzen (2020), then sequentially predicts the answers to subproblems, and finally generating the final output. Compositionality in NLP is closely related with multi-hop QA problems with compositional questions where the answers from sub-questions are needed to answer the main question Yang et al. (2018); Ho et al. (2020); Trivedi et al. (2022). Qiu et al. (2022) have shown how a model with compositional latent structure improves large language models’ performance on compositional generalization tasks through synthethic data augmentation. CoI is most related to Aksu et al. (2023) as they work on dealing with compositional tasks for dialogue systems. However, their definition of “compositional task” is different from ours as they do not require the output of one subtask is shown to the next subtask. Meanwhile, in our CoI, the outputs from the previous subtasks become the next subtask inputs.
6 Conclusion and Future Work
In this work, we propose a new task called Chain-of-Instructions and develop a dataset for building models to solve the task. We introduce an automatic pipeline on how to build our dataset and demonstrate the usefulness of our CoI-tuned models on the tasks of the generated dataset and downstream tasks. Since human language is complex and an instruction may actually be composed of subtasks, it is important to have a model that can deal with compositional instructions, especially as we show that models fine-tuned only on single instructions never outperform the CoI-tuned models on multi-instruction tasks. For future work, we consider looking into instruction decomposition in addition to the instruction composition problem.
Limitations
Our second and third outputs are generated by a large language model, GPT-3.5111111https://platform.openai.com/docs/models/gpt-3-5-turbo, instead of humans. Thus, a 100% correctness of these outputs is not guaranteed. However, through human evaluation on a randomly selected 100 samples, we found that around 75% of them have correct second outputs. For the 3-instruction composition, 60% out of 100 randomly sampled instances have valid responses.
While our study only covers a smaller variant of LLMs (the 7B model), our findings demonstrates that instructing the model with compositional guidelines significantly enhances its ability to understand and apply compositional principles, as CoI models outperformed non-fine-tuned baselines. Although we have not conducted experiments with larger LLMs, employing larger models entail higher training costs. Nevertheless, our findings suggest that even fine-tuning smaller LLMs can impart compositional knowledge to the model, resulting in improved generalizability for both downstream tasks and single-instruction problem.
Appendix A Appendix
A.1 CoI Dataset Statistics
@@ A.2 Prompting Details
A.2.1 Summarizing and generalizing a single instruction
@@
A.2.2 Checking compositionality of two instructions and obtaining the second output
@@
A.2.3 Making the input type of the second instruction consistent with the output type of the first instruction
@@
A.2.4 Separating outputs for the subtasks from Baseline Outputs
@@ A.2.5 Creating concise-CoI
A.3 Fine-tuning Details
@@
A.4 BIG-Bench Hard Examples
@@ A.5 WikiLingua Examples
A.6 Data Generation Details
@@
Single Instruction from Sup-NatIns Wang et al. (2022)
@@
A.7 Heuristics for Validity Check
@@ A.9 Full results
A.10 Full Intermediate Results

### v3

@@ Results
Several NLP work have investigated the capability of Transformer model on compositional problems including algorithm and math problems (Dziri et al. 2023), compositional semantics (Drozdov et al. 2022), and multi-hop question-answering (QA) tasks (Trivedi et al. 2022). Dziri et al. (2023) highlight how Transformer models often struggle with compositional mathematics computation or program executions (Nye et al. 2022; Saparov and He 2022). Drozdov et al. (2022) introduce a new prompting method which first decomposes the compositional questions or sentences (Keysers et al. 2019; Kim and Linzen 2020), then sequentially predicts the answers to subproblems, and finally generating the final output. Compositionality in NLP is closely related with multi-hop QA problems with compositional questions where the answers from sub-questions are needed to answer the main question (Yang et al. 2018; Ho et al. 2020; Trivedi et al. 2022). Qiu et al. (2022) have shown how a model with compositional latent structure improves large language models’ performance on compositional generalization tasks through synthethic data augmentation. CoI is most related to Aksu et al. (2023) as they work on dealing with compositional tasks for dialogue systems. However, their definition of “compositional task” is different from ours as they do not require the output of one subtask is shown to the next subtask. Meanwhile, in our CoI, the outputs from the previous subtasks become the next subtask inputs.
Conclusion and Future Work
In this work, we propose a new task called Chain-of-Instructions and develop a dataset for building models to solve the task. We introduce an automatic pipeline on how to build our dataset and demonstrate the usefulness of our CoI-tuned models on the tasks of the generated dataset and downstream tasks. Since human language is complex and an instruction may actually be composed of subtasks, it is important to have a model that can deal with compositional instructions, especially as we show that models fine-tuned only on single instructions never outperform the CoI-tuned models on multi-instruction tasks. For future work, we consider looking into instruction decomposition in addition to the instruction composition problem. We also recommend trying out more tasks to be composed besides those from SuperNaturalInstruction.
Acknowledgements
We would like to thank members of the MinnesotaNLP lab for their feedback and intellectual support.
Appendix
CoI Dataset Statistics
@@ Acknowledgements
Summarizing and generalizing a single instruction
@@
Checking compositionality of two instructions and obtaining the second output
@@
Making the input type of the second instruction consistent with the output type of the first instruction
@@
Separating outputs for the subtasks from Baseline Outputs
@@
LLM as a Preference Judge
The following is the prompt for LLM as a preference judge:
Given the following instruction and input and ground truth output, which generated output follows the instruction more closely?
Answer only with “A ”if you prefer generated output A, “B” if you prefer generated output B, “None” if you prefer none.
Instruction: instruction Input: input Ground truth output: ground truth output
Generated output A: output from model A Generated output B: output from model B
Fine-tuning Details
@@
BIG-Bench Hard Examples
@@
Data Generation Details
@@
Single Instruction from Sup-NatIns (Wang et al. 2022)
@@
Heuristics for Validity Check
@@
Full Intermediate Results